# Diffusion Suction Grasping with Large-Scale Parcel Dataset

### v1

@@ Abstract
While recent advances in object suction grasping have shown remarkable progress, significant challenges persist particularly in cluttered and complex parcel handling scenarios. Two fundamental limitations hinder current approaches: (1) the lack of a comprehensive suction grasp dataset tailored for parcel manipulation tasks, and (2) insufficient adaptability to diverse object characteristics including size variations, geometric complexity, and textural diversity. To address these challenges, we present Parcel-Suction-Dataset, a large-scale synthetic dataset containing 25 thousand cluttered scenes with 410 million precision-annotated suction grasp poses. This dataset is generated through our novel geometric sampling algorithm that enables efficient generation of optimal suction grasps incorporating both physical constraints and material properties. We further propose Diffusion-Suction, an innovative framework that reformulates suction grasp prediction as a conditional generation task through denoising diffusion probabilistic models. Our method iteratively refines random noise into suction grasp score maps through visual-conditioned guidance from point cloud observations, effectively learning spatial point-wise affordances from our synthetic dataset. Extensive experiments demonstrate that the simple yet efficient Diffusion-Suction achieves new state-of-the-art performance compared to previous models on both Parcel-Suction-Dataset and the public SuctionNet-1Billion benchmark.
@@ I Introduction
With the growing popularity of the e-commerce industry, logistics companies are facing a heavy burden in processing and shipping the rapidly increasing volume of parcels [1]. At present, traditional manual processing can no longer cope with such increasing parcels [2]. However, due to the disorderly stacked and diverse types of express packages, processing and shipping parcels face technical challenges. In particular, accurate and quick object grasping has become a fundamental issue in the effective implementation of vision-based intelligent parcels processing system.
@@
Due to the simplicity and reliability, suction grasping is widely used in real-world tasks in many industrial and daily scenes, among the different grasping techniques [3]. Despite the success of recent object suction grasping methods, challenges still exist and need to be addressed especially in the context of the cluttered and complex parcel scenes. Two major challenges persist in obtaining robust and accurate suction grasping for parcel object. The first challenge lies in the absence of a universal and efficient suction grasping datasets for parcel scenes. Zeng et al. [4] solely relied on manual experience to annotate the suctionable area and previous works [5] proposed a synthetic suction dataset collection pipeline for cluttered environment. These dataset lack a general focus on the characteristics of parcel objects and a substantial amount of flat object data features. Consequently, models trained on such datasets cannot directly generalize to complex parcel scenes. Moreover, most grasp datasets samples cannot adequately account for the characteristics of large-area and severe-obstruction of objects, resulting in failed suction grasping in real-world scene.
@@
Figure 1: This work addresses two major problems in suction grasping task. We propose a novel pipeline which evaluate suction grasping from different perspectives to obtain annotation labels. We propose a novel framework to predict suction grasping pose by reformulating it as an iterative diffusion-denoising process.
@@
Secondly, suction grasping prediction is a challenging task because it involves finding feasible suction grasping poses given the wide range of object sizes, shapes and textures. Current methods directly formulate this problem as a regression problem. Many object 6D pose estimation models [6, 7, 8, 9] project the pre-defined suction configuration onto the scene. However, these approaches lack generalization for new objects. Alternatively, discriminative approaches [4, 10, 5] directly generate suction grasping poses from point clouds or RGB-D images. Nonetheless, these approaches are frequently limited by specialized network architectures and diverse training strategies to generate reliable predictions.
@@
In this study, we introduce the Parcel-Suction-Dataset and Diffusion-Suction for suction grasping prediction in cluttered scenes, as shown in Fig. 1. We propose a novel Self-Parcel-Suction-Labeling framework which can automatically generate general and diverse dataset information in parcel stacked scenes. To the best of our knowledge, this comprehensive dataset is the first large-scale synthetic suction grasping dataset in parcel stacked scenes, called Parcel-Suction-Dataset. Specifically, we propose an image to 3D parcel asset pipeline which can comprehensively capture both geometry and appearance information to obtain 3D high-quality assets. Moreover, we use Bullet and Blender platform to create random unstructured parcel scenes and generate realistic synthetic rendering information, such as RGB images, segmentation images and depth maps. We introduce a suction grasping candidate evaluation algorithm that calculate suction cup gripper seal score, wrench score, collision score and visibility score without requiring labor-intensive processes.
@@
Moreover, we propose a novel Diffusion-Suction framework to predict feasible and accurate suction grasping poses by reformulating the problem as an iterative diffusion-denoising refinement process with 3D visual-condition guided. Diffusion-Suction is decoupled into the point clouds encoder and suction grasping score decoder. It uses advanced point clouds feature extraction network PointNet++ to extract global information as condition guidance. Subsequently, we propose a novel and light-weighted Pointcloud Conditioned Denoising Block which conduct Attention on the hierarchical channel and spatial information to emphasize the significant features. The point clouds encoder is executed in a single time and the diffusion process is performed using a light-weighted decoder head. With this simple and efficient structure, Diffusion-Suction significantly reduces the inference computational overhead. During training stage, gaussian noise controlled by a variance schedule [11] is iteratively added to ground truth to obtain noisy maps. At the inference stage, Diffusion-Suction generates reliable suction grasping score by reversing the learned diffusion process. To the best of our knowledge, this is the first work introducing the diffusion model into suction grasping prediction task.
@@
We evaluate our proposed method using the Parcel-Suction-Dataset and public SuctionNet-1Billion [3] benchmarks. Our experimental results demonstrate that Diffusion-Suction significantly outperforms state-of-the-art methods. We present extensive ablation study to further reveal the effectiveness and properties of Diffusion-Suction on Parcel-Suction-Dataset. In summary, the main contributions of our work are:
@@
  • We propose a novel Self-Parcel-Suction-Labeling framework to generate general and diverse large-scale Parcel-Suction-Dataset in parcel stacked scenes.
@@
  • We formulate suction grasping prediction task as a generative denoising process, which is the first study to apply the diffusion model to this task.
@@
  • We propose a novel Diffusion-Suction framework to predict feasible and accurate suction grasping pose with 3D visual-condition guided.
@@ II-A Suction Dataset
Zeng et al. [4] conducted a manual annotated real-world dataset that contains cluttered scenes. Its disadvantage is that it relies solely on manual experience annotation for the suctionable and non-suctionable area, making the annotation process highly time-consuming and expensive. In addition, manual annotation introduces significant potential errors. Jiang et al. [12] proposed a reachability evaluation metric to avoid problems when a high grasp quality configuration is unable to reach the target. Shao et al. [13] conducted a self-supervised learning robotic bin-picking system to generate training data and completed training within a few hours. But it has only been validated on cylindrical objects, limiting the application of this method. Suctionnet [3] conducted a real word dataset which utilizes a new physical model to analytically evaluate seal and wrench formation. The annotation error depends on the accuracy of pose estimation in real scenes. CoAS-Net. [5] proposed a suction dataset collection synthetic pipeline that contains collision-free annotation and uses the newest simulator Issac Sim to reduce the domain gap. Sim-Suction [14] proposed a large-scale synthetic dataset for cluttered environments, which uses object-aware suction grasp sampling and considers dynamic interactions in unstructured environments. However, there is currently a lack of readily available data in parcel fields to explore and improve suction grasping. We propose Self-Parcel-Suction-Labeling framework to build large-scale parcel suction dataset which includes a variety of parcel objects acquired in the real world. In addition, our dataset is specifically designed for parcel scenes, which is severely stacked and contains a substantial amount of flat object data features.
Figure 2: Overview of Self-Parcel-Suction-Labeling pipeline. Firstly, utilize the image to 3D parcel asset model to generate 3D high-quality assets with geometry and appearance information. Next, create random unstructured parcels scenes with the Bullet and Blender simulator platform. Finally, evaluate suction grasping candidates from four different perspectives to obtain accurate annotation labels.
@@ II-B Suction-based Grasping
Hernandez et al. [6] estimated the 6D pose of the object [7, 8, 9] and projected the pre-defined suction configuration onto the objects in the scene. However, such methods lack generalization for new objects. Zeng et al. [4] design a single FCN architecture trained on a human-labeled dataset, learning pixel-wise suction affordances. Suctionnet [3] proposed pixel-wise suction scores prediction network which predicts seal score heatmap and center score heatmap separately. Sim-Suction [14] proposed a object-aware affordance network which directly inputs text prompts to obtain identify regions of interest grasp object and outputs point-wise suction probability. Zhang et al. [15] propose the suction reliability matrix and suction region prediction model. However, it has not been tested for effectiveness in multi object scenes. CoAS-Net. [5] utilized context aware suction network trained with a synthetic dataset. Rui Cao et al. [10] proposed an uncertainty-aware and multi-stage framework which exploit both aleatoric and epistemic uncertainties. These methods essentially belong to discriminative approaches and generative approaches have not yet been explored in terms of suction grasping task. Diffusion models [16] demonstrate remarkable performance in text-to-image generation task and surpass previous generative models, such as Generative Adversarial Networks (GAN) [17] or VAE [18]. We propose to utilize the diffusion model to denoise random noise into suction grasping score map with the guidance of input point clouds visual-condition, instead of adopting it as a normal regression head. To the best of our knowledge, this is the first work introducing the diffusion model into suction grasping prediction task.
@@ III Large-Scale Parcel Suction Dataset
As mentioned earlier, parcel scene has not been widely covered in existing datasets. Therefore, we propose a Self-Parcel-Suction-Labeling (SPSL) framework which contains diverse information and constructs Parcel-Suction-Dataset as shown in Fig. 2. This section will provide more detailed information about the SPSL.
@@ III-A Randomized Scene Generation
Acquiring 3D high-quality assets is the first step in generating scenes for dataset construction. However, there is currently a lack of available 3D parcel assets. The 3D parcel assets have the characteristics of privacy, complexity, and diversity. Using traditional laser scanning instruments to obtain 3D assets through scanning cannot effectively solve the above problems and is time-consuming [19]. To address this issue, we propose a pipeline that converts images to 3D parcel assets. Specifically, we collect multiview images in real parcel stacked scenes as prompts for the latest 3D generation method TRELLIS [20] which can comprehensively capture both structural (geometry) and textural (appearance) information. We have generated a total of 113 high-quality parcel assets and clustered them in three categories according to the Chamfer Distance and Normal Consistency metric: rectangular, planar, and cylindrical.
We create random unstructured parcel scenes using the previously generated assets. In the real world, parcels are stacked randomly into a pile. To simulate this real state, parcels are randomly dropped into a bin with arbitrary poses on the Bullet simulator platform [21] which can simulate the real dynamical collision. Moreover, we use Blender platform [22] to generate realistic and accurate synthetic rendering information, such as RGB images, segmentation images and depth maps. To reduce the domain gap between the synthetic and real domains, we randomize several aspects of the simulation.
  • Randomly select objects from the parcel dataset.
@@ III-B Suction Pose Annotation
At this stage, we will describe how to compute the suction grasping score, meaning that the higher the score, the easier it is to suction grasp in real-world scene. We assign them four continuous scores, namely seal score, wrench score, collision score and visibility score. We define the final metric as a product of the sub-evaluation score $S={S_{seal}}\times{S_{wrench}}\times{S_{collision}}\times{S_{visibility}}$.
Seal Score Evaluation. The seal score is used to determine whether the suction cup can maintain a vacuum state when suction grasping is performed in a specific pose. The higher the surface flatness of the object, the higher the seal score. Similar to Cao et al [3], we leverage the compliant suction contact model [4] to evaluate this process, which models the suction cup as a quasi-static spring system. Specifically, we select a suction cup gripper with a 10 mm radius suction cup and sample dense vertex at the suction cup $\{{v_{1}},{v_{2}},...,{v_{n}}\}$. Then we transform the suction cup model to the object surface along a specific approach vector to obtain object projection surface vertex $\{\widetilde{{v_{1}}},\widetilde{{v_{2}}},...,\widetilde{{v_{n}}}\}$. We use the distance variation between sampling points to evaluate the seal score, which can be defined as:
@@
where the i-th sampling point distance variation ${r_{i}}=min(1,\frac{{\widetilde{{l_{i}}}-{l_{i}}}}{{{l_{i}}}})$ and the i-th distance between adjacent sampling point ${l_{i}}={v_{i}}-{v_{i+1}}$. The weight of objects is not considered during seal score evaluation process. Note that we control the sampling points based on the size changes of the object, in order to achieve an optimal balance between accuracy and efficiency.
@@
where $\mu$ is the friction constant, ${V_{f}}$ is suction grasping force, ${f_{f}}=({f_{x}},{f_{y}})$ is frictional force, ${{\tau_{z}}}$ is torsional friction, ${\tau_{e}}=({\tau_{x}},{\tau_{y}})$ is elastic restoring torque, $r$ is the radius suction cup and $k$ is a material-related constant. Similar to Cao et al [3], we have simplified these conditions by reserving only the material condition to evaluate this process. In the real world, suction grasp objects from the side often leads to failure, and humans are accustomed to suction grasp objects from the top. However, for planar category objects in parcel dataset as mentioned earlier, due to the short distance from the suction grasping point to the center of mass of the object when suction grasping from the side, this evaluation algorithm obtains a high wrench score. Therefore, we propose a correction strategy that uses the angle between the suction grasping normal vector and gravity to correct. Based on previous assumption, the wrench score can be defined as:
${S_{wrench}}=(1-min(1,\left|{{\tau_{e}}})\right|/{\tau_{thr}})(1-a/\pi)$ (6)
@@
Collision Score Evaluation. The collision score is used to determine whether the suction cup collides with objects in a specific pose in the scene. We check whether any points in the scene appear inside the mesh of the suction cup gripper. If there are points in the gripper, ${S_{{\rm{collision}}}}$ is 0, and vice versa. To balance between computational cost and accuracy, we set up a workspace around the suction cup gripper for collision detection.
Visibility Score Evaluation. The visibility score is used to reflect the obstruction of the object in the scene. In highly occluded parcel scenes, some parcels are severely occluded and we are not interested in these parcel instances which cannot be captured at the bottom. Moreover, when these parcels are grabbed, it will seriously affect other objects in the scene, causing grabbing failure. Specifically, given an occluded scene image where parcels are stacked randomly into a pile, we calculate the pixel area size of each parcel $\widetilde{{p_{o}}}$. Then, we iteratively remove other objects in the scene and calculate the unobstructed pixel area size of each object ${p_{o}}$. Visibility score can be defined as ${S_{{\rm{visibility}}}}=\widetilde{{p_{o}}}/{p_{o}}$.

### v2

@@ Abstract
While recent advances in suction grasping have shown remarkable progress, significant challenges persist particularly in cluttered and complex parcel handling scenarios. Current approaches are limited by (1) the lack of comprehensive parcel-specific suction grasp datasets and (2) poor adaptability to diverse object properties, including size, geometry, and texture. We address these challenges through two main contributions. Firstly, we introduce the Parcel-Suction-Dataset, a large-scale synthetic dataset containing 25 thousand cluttered scenes with 410 million precision-annotated suction grasp poses, generated via our novel geometric sampling algorithm. Secondly, we propose Diffusion-Suction, a framework that innovatively reformulates suction grasp prediction as a conditional generation task using denoising diffusion probabilistic models. Our method iteratively refines random noise into suction grasping score through visual-conditioned guidance from point cloud observations, effectively learning spatial point-wise affordances from our synthetic dataset. Extensive experiments demonstrate that the simple yet efficient Diffusion-Suction achieves new state-of-the-art performance compared to previous models on both Parcel-Suction-Dataset and the public SuctionNet-1Billion benchmark. This work provides a robust foundation for advancing automated parcel handling systems in real-world applications.
@@ I Introduction
With the growing popularity of the e-commerce industry, logistics companies are facing an increasing burden in processing and shipping the rapidly increasing volume of parcels [1]. At present, traditional manual processing are no longer capable of handling such a surge in parcel volume [2]. However, due to the disorderly stacked and diverse types of express packages, processing and shipping parcels present significant technical challenges. In particular, accurate and rapid object grasping has become a fundamental issue in the effective implementation of vision-based intelligent parcels processing systems.
@@
Due to its simplicity and reliability, suction grasping is widely used in real-world tasks across various industrial and daily scenes, among the different grasping techniques [3]. Despite the success of recent object suction grasping methods, challenges still exist and need to be addressed especially in the context of the cluttered and complex parcel scenes. Two major challenges persist in obtaining robust and accurate suction grasping for parcel objects. The first challenge lies in the absence of a universal and efficient suction grasping datasets for parcel scenes. Zeng et al. [4] solely relied on manual experience to annotate the suctionable area and previous works [5] proposed a synthetic suction dataset collection pipeline for cluttered environment. However, these datasets lack a general focus on the characteristics of parcel objects and a substantial amount of flat object data features. Consequently, models trained on such datasets cannot directly generalize to complex parcel scenes. Moreover, most grasp dataset samples cannot adequately account for the characteristics of large-area and severe-obstruction of objects, leading to failed suction grasping in real-world scenes.
@@
Figure 1: This work addresses two major challenges in the suction grasping task. We propose a novel pipeline that evaluates suction grasping from different perspectives to obtain annotation labels. We propose a novel framework to predict suction grasping poses by reformulating the task as an iterative diffusion-denoising process.
@@
Secondly, suction grasping prediction is a challenging task because it involves finding feasible suction grasping poses given the wide range of object sizes, shapes, and textures. Many object 6D pose estimation models [6, 7, 8] project the pre-defined suction configurations onto the scene. However, these approaches lack generalization capabilities for new objects. Alternatively, discriminative approaches [4, 9, 5] directly generate suction grasping poses from point clouds or RGB-D images. Nonetheless, these approaches are often limited by specialized network architectures and diverse training strategies to generate reliable predictions.
@@
In this study, we introduce the Parcel-Suction-Dataset and Diffusion-Suction for suction grasping prediction in cluttered scenes, as shown in Fig. 1. We propose a novel Self-Parcel-Suction-Labeling framework, which can automatically generate general and diverse dataset information for parcel-stacked scenes. To the best of our knowledge, this comprehensive dataset is the first large-scale synthetic suction grasping dataset specially designed for parcel stacked scenes, named Parcel-Suction-Dataset. Specifically, we propose an image to 3D parcel asset pipeline that can comprehensively capture both geometry and appearance information to obtain high-quality 3D assets. Moreover, we utilize Bullet and Blender platforms to create random unstructured parcel scenes and generate realistic synthetic rendering information, such as RGB images, segmentation images, and depth maps. We introduce a suction grasping candidate evaluation algorithm that calculates suction cup gripper’s seal score, wrench score, collision score and visibility score without requiring labor-intensive processes.
@@
Moreover, we propose a novel Diffusion-Suction framework to predict feasible and accurate suction grasping poses by reformulating the problem as an iterative diffusion-denoising refinement process guided by 3D visual conditions. Diffusion-Suction is decoupled into the point cloud encoder and suction grasping score decoder. It utilizes the advanced point cloud feature extraction network to extract global information as condition guidance. Subsequently, we propose a novel and lightweight Pointcloud Conditioned Denoising Block, which conducts attention on the hierarchical channel and spatial information to emphasize the significant features. The point cloud encoder is executed once and the diffusion process is performed using a lightweight decoder head. With this simple and efficient structure, Diffusion-Suction significantly reduces the computational overhead during inference. During the training stage, Gaussian noise controlled by a variance schedule [10] is iteratively added to the ground truth to obtain noisy maps. At the inference stage, Diffusion-Suction generates reliable suction grasping scores by reversing the learned diffusion process. To the best of our knowledge, this is the first work introducing the diffusion model into suction grasping prediction task.
@@
We evaluate our proposed method using the Parcel-Suction-Dataset and the public SuctionNet-1Billion [3] benchmark. Our experimental results demonstrate that Diffusion-Suction significantly outperforms state-of-the-art methods. We present an extensive ablation study to further reveal the effectiveness and properties of Diffusion-Suction on the Parcel-Suction-Dataset. In summary, the main contributions of our work are as follows:
@@
  • We propose a novel Self-Parcel-Suction-Labeling framework to generate general and diverse large-scale Parcel-Suction-Dataset for parcel-stacked scenes.
@@
  • We formulate the suction grasping prediction task as a generative denoising process, which is the first study to apply the diffusion model to this task.
@@
  • We propose a novel Diffusion-Suction framework to predict feasible and accurate suction grasping poses guided by 3D visual conditions.
@@ II-A Suction Dataset
Zeng et al. [4] conducted a manual annotated dataset containing cluttered scenes. However, it relies solely on manual experience for annotating the suctionable and non-suctionable areas, making the annotation process highly time-consuming and expensive. Shao et al. [11] conducted a self-supervised learning robotic bin-picking system to generate training data and completed training within a few hours. However, it has only been validated on cylindrical objects, limiting the application of this method. Suctionnet [3] conducted a real word dataset which utilizes a new physical model to analytically evaluate seal and wrench formation. The annotation error depends on the accuracy of pose estimation in real scenes. Sim-Suction [12] proposed a large-scale synthetic dataset for cluttered environments, which uses object-aware suction grasp sampling. However, there is currently a lack of readily available data in the parcel fields to explore and improve suction grasping. We propose a Self-Parcel-Suction-Labeling framework to build a large-scale parcel suction dataset which includes a variety of parcel objects collected from the real world. Furthermore, our dataset is specifically designed for parcel scenes, featuring densely stacked packages and a significant amount of flat objects.
Figure 2: Overview of Self-Parcel-Suction-Labeling pipeline. Firstly, utilize the image prompts to generate 3D parcel asset model to generate high-quality 3D assets with geometry and appearance information. Next, create random unstructured parcel scenes with the Bullet and Blender simulator platform. Finally, evaluate candidate suction grasps from four different perspectives to obtain accurate annotation labels.
@@ II-B Suction-based Grasping
Hernandez et al. [6] estimated the 6D pose of the object [7, 13, 14, 15, 8] and projected the pre-defined suction configuration onto the objects in the scene. However, such methods lack generalization for new objects. Zeng et al. [4] designed a single FCN architecture trained on a human-labeled dataset, learning pixel-wise suction affordances. Suctionnet [3] proposed a pixel-wise suction scores prediction network which predicts seal score heatmap and center score heatmap separately. Sim-Suction [12] proposed a object-aware affordance network which directly inputs text prompts to obtain identify regions of interest grasp object and outputs point-wise suction probability. Zhang et al. [16] proposed the suction reliability matrix and suction region prediction model. However, these methods have not been tested for their effectiveness in multi-object scenes. CoAS-Net [5] proposed a context aware suction network trained on a synthetic dataset. Rui Cao et al. [9] proposed an uncertainty-aware and multi-stage framework which exploit both aleatoric and epistemic uncertainties. These methods essentially belong to discriminative approaches, while generative approaches have not yet been explored in terms of suction grasping task. Diffusion models [17] demonstrate remarkable performance in text-to-image generation task and surpass previous generative models, such as Generative Adversarial Networks (GAN) [18] or VAE [19]. We propose to utilize the diffusion model to generate a suction grasping score map from random noise with the guidance of the visual information of the input point cloud, instead of adopting it as a normal regression head. To the best of our knowledge, this is the first work introducing the diffusion model into the suction grasping prediction task.
@@ III Large-Scale Parcel Suction Dataset
As mentioned earlier, the parcel scene has not been widely covered in existing datasets. Therefore, we propose a Self-Parcel-Suction-Labeling (SPSL) framework incorporating diverse information and construct the Parcel-Suction-Dataset as shown in Fig. 2. This section will provide more detailed information about the SPSL.
@@ III-A Randomized Scene Generation
Acquiring high-quality 3D assets is the first step in generating scenes for dataset construction. However, there is currently a lack of available 3D parcel assets. The 3D parcel assets have the characteristics of privacy, complexity, and diversity. Using traditional laser scanning instruments to obtain 3D assets through scanning cannot effectively solve the above problems and is time-consuming. To address this issue, we propose an efficient and effective pipeline that converts images to 3D parcel assets. Specifically, we collect multiview images in real parcel stacked scenes as prompts for the latest 3D generation method TRELLIS [20] which can comprehensively capture both structural (geometry) and textural (appearance) information. We have generated a total of 113 high-quality parcel assets and clustered them into three categories according to the Chamfer Distance and Normal Consistency metric: rectangular, planar, and cylindrical.
Using the previously generated assets, we create random unstructured parcel scenes. In the real world, parcels of various shapes and sizes are stacked randomly and disorderly into a pile. To simulate this real-world situation, we utilize the Bullet simulator platform, which is capable of simulating real dynamical collisions. Parcels are then randomly dropped into a bin with arbitrary poses on this platform. Furthermore, the Blender platform is employed to generate realistic and accurate synthetic rendering information, including RGB images, segmentation images, and depth maps. To reduce the domain gap between the synthetic and real domains, we randomize several crucial aspects of the simulation.
@@ III-B Suction Pose Annotation
At this stage, we will describe how to compute the suction grasping score, and the higher the score is, the easier it will be to perform suction grasping in real-world scenes. We assign four continuous scores for each suction grasping situation, namely seal score, wrench score, collision score and visibility score. We define the final metric as the product of the sub-evaluation scores $S={S_{seal}}\times{S_{wrench}}\times{S_{collision}}\times{S_{visibility}}$.
Seal Score Evaluation. The seal score is used to determine whether the suction cup can maintain a vacuum state when suction grasping is performed in a specific pose. The higher the surface flatness of the object, the higher the seal score. Similar to Cao et al [3], we leverage the compliant suction contact model [4] to evaluate this process, which models the suction cup as a quasi-static spring system. Specifically, we sample dense vertices on the suction cup surface, denoted as $\{{v_{1}},{v_{2}},...,{v_{n}}\}$. Then we transform the suction cup model to the object surface along a specific approach vector to obtain object projection surface vertices $\{\widetilde{{v_{1}}},\widetilde{{v_{2}}},...,\widetilde{{v_{n}}}\}$. We use the distance variation between sampling points to evaluate the seal score, which can be defined as:
@@
where the $i$-th sampling point distance variation ${r_{i}}=\min\left(1,\frac{{\widetilde{{l_{i}}}-{l_{i}}}}{{{l_{i}}}}\right)$. Here, $l_{i}$ is the distance between two adjacent sampling points $v_{i}$ and $v_{i+1}$ on the surface of the suction cup before the transformation of the suction cup model to the object surface. $\widetilde{l_{i}}$ represents the corresponding distance between the projected points of $\widetilde{{v_{i}}}$ and $\widetilde{{v_{i+1}}}$ on the object surface after the transformation. The weight of objects is not considered during seal score evaluation process. Note that we control the sampling points based on the size changes of the object, in order to achieve an optimal balance between accuracy and efficiency.
@@
where $\mu$ is the friction constant, ${V_{f}}$ is suction grasping force, ${f_{f}}=({f_{x}},{f_{y}})$ is frictional force, ${{\tau_{z}}}$ is torsional friction, ${\tau_{e}}=({\tau_{x}},{\tau_{y}})$ is elastic restoring torque, $r$ is the radius suction cup and $k$ is a material-related constant. Similar to Cao et al. [3], we have simplified these conditions by retaining only the material condition to evaluate this process. In the real world, suction grasping objects from the side often leads to failure, and humans are accustomed to suction grasping objects from the top. However, when suction grasping these objects from the side, the distance from the suction grasping point to the center of mass of the object is short. As a result, this evaluation algorithm obtains a high wrench score. Therefore, we propose a correction strategy that uses the angle between the suction grasping normal vector and gravity to correct the wrench score. Based on previous assumption, the wrench score can be defined as:
${S_{wrench}}=(1-min(1,\left|{{\tau_{e}}})\right/{\tau_{thr}})(1-a/\pi)$ (6)
@@
Collision Score Evaluation. The collision score is used to determine whether the suction cup collides with objects in a specific pose in the scene. We check whether any points in the scene appear inside the mesh of the suction cup gripper. If there are points detected within the gripper, ${S_{{\rm{collision}}}}$ is set to 0, and vice versa. To balance between computational cost and accuracy, we set up a workspace around the suction cup gripper for collision detection.
Visibility Score Evaluation. The visibility score is used to quantitatively reflect the degree of obstruction of the object in the scene. In highly occluded parcel scenes, some parcels are severely occluded and we are not interested in these parcel instances which cannot be captured at the bottom. Moreover, when these parcels are grabbed, it will seriously affect other objects in the scene, causing grabbing failure. Specifically, given an occluded scene image where parcels are stacked randomly into a pile, we calculate the pixel area size of each parcel, denoted as $\widetilde{{p_{o}}}$. Then, we iteratively remove other objects in the scene and calculate the unobstructed pixel area size of each object ${p_{o}}$. Visibility score can be defined as ${S_{{\rm{visibility}}}}=\widetilde{{p_{o}}}/{p_{o}}$.